Given that the sum of displacement in the tth  second and (t+1)th second is 100cm. Find the velocity after t seconds   

Grade:12

1 Answers

SOURAV MISHRA
10 years ago

St = u + a*(2t - 1)/2 and St + 1 = u + a*(2t + 1)/2 

we have St + St + 1 = 2u + 2at = 2(u + at) = 2v = 100

v = 50 cm/s.

therefore the velocity after t seconds is 50 cm/s.